# A Decision-Language Model (DLM) for Dynamic Restless Multi-Armed Bandit Tasks in Public Health

### 22 Feb 2024 v1

@@ 5.4 DLM Reflection Example
We visualize for two prompts, two candidate reward distributions and the selected best index in Figure 3. We show 4 relevant feature categories of age, income, education, and phone owners for both prompts. We observe that the DLM selection is able to reasonably interpret commands, such as prioritizing older ages, and select the distribution that best achieves the given prompt. However, we note that for more ambiguous criteria, such as “technologically challenged”, DLM selection may be more opaque; while the base reward for this prompt considers only income and phone ownership, additional features such as education level may also be considered for selection. In these cases, we hypothesize that additional human feedback may be required to better disambiguate prompts for improved reflection. We provide the full reward reflection text for the below examples in Appendix E.
Figure 3: DLM reflection distribution and selection visualization. DLM uses outcome distributions from trained policy simulations to refine proposed rewards. Above, we show the selected best candidate reward for two prompts. DLM is able to reasonably interpret given prompts using distributions across feature groups; however, for more ambiguous prompts (“technologically challenged”), the selection criteria may become more opaque, potentially requiring additional human feedback.
@@ 5.5 Exceeding Base in Critical Low Income
We find that empirically for the “Critical Low Income” task the LLM is actually able to improve over the base reward. We hypothesize here that the DLM model may propose a reward that is of much larger scaling than the original reward. For a representative example, we show a sample reward generated by the LLM for the Critical Low Income prompt:
@@
state + if_(state) * 10 * if_(agent_feats.no_or_lowest_income or agent_feats.second_lowest_income or agent_feats.third_lowest_income),
As the results reported in Figure 2 are taken in the low finite-sample regime, where there is not much data to train from, sample efficiency becomes extremely important. In this case, because the reward function generated by the LLM gives an order of magnitude higher weighting to specific arms, it seems likely that the underlying policy gradient algorithm PPO gains a large amount of sample efficiency in our setting. It is important to note that the improvement that the LLM yields over the base reward will shrink as the number of samples taken during training goes to infinity, as by definition the base reward defines the optimal policy. However, the highly shaped reward given by the LLM in this setting helps PPO converge faster for this task. It is also important to note that the base reward may not be the most sample efficient reward; other more complex rewards relying on correlated features or with different reward scaling are likely more efficient. Thus, this opens the possibility of achieving potential superhuman performance on our task in the future.
@@ 6 Conclusion
We propose a Decision-Language Model for resource allocation tasks in public health. The proposed model can take language prompts describing complex, potentially ambiguous policy outcome goals as input, and generate downstream control policies specialized for such goals. The proposed LLM-based approach can accurately capture the referenced prompt features and design effective reward functions in the form of code. Critically, our algorithms enable self-iteration on proposed reward functions in a fully automated fashion. By grounding LLM reward proposals with outcome distributions from RMAB simulations, we demonstrate the ability for the proposed DLM model to achieve near-base performance across a variety of policy preference tasks.
@@
Consent is obtained from every beneficiary enrolling in the ARMMAN’s mobile health program (see also Appendix A). For future deployment of the proposed algorithm, one could take extra steps to avoid unsafe policies, including: (1) examining per beneficiary policy outcomes and performing spot-checks on trajectories; (2) apply the proposed algorithm to a subgroup of the population; (3) provide additional context to LLM after observing the impact on different features. We discuss measures to address fairness in Appendix B.
Appendix
@@ Appendix D Provided Tasks
We show provided tasks in Table 1. We note that we evaluate on a wide variety of tasks within the real of resource allocation for maternal health. In particular, we test across feature characteristics such as age, income, spoken languages (an important characteristic indicating ethnicity in the Indian subcontienent), education, and other relevant features that may be considered in real-world allocation tasks. We provide these tasks directly as context to the LLM, in addition to relevant feature characteristics and their corresponding descriptions. Per seed, we fix the randomness in the sampled set of 48 arms from ARMMAN’s dataset, the deep RL neural network initialization, batch sampling from the dataset, and the sampling of states for each arm during evaluation. However please note that we cannot control the randomness of the Gemini LLM generation as we only have access through the Google API interface.
Table 1: Full prompt list
Label Task Command Base Function No Action Base Human Default DLM (No Reflection) DLM (Reflection)
@@
Combine the distributions of ‘age’ and ‘education’ to heavily give precedence to low income impoverished youth.
@@
Infer technical challenges in reaching the phone that could indicate ‘at-risk’ beneficiaries and give preference.
@@
Appendix E Sample Reward Reflection Outputs
@@
E.1 Task: Older Age Bias
@@ 35. Income bracket -1 (no income) - Binary
E.2 Task: Technologically Challenged
@@
Appendix F Generated Reward Functions
Here we give all generated reward functions seen in Table 1 from the LLM. We first give the task description, which is fed to the LLM, the corresponding ground truth base reward, and then the generated LLM rewards in order of seed. Note that we have processed the original raw indices into feature labels for interpretability.
F.1 Focus on the most elderly by age distribution.
@@ Appendix F Generated Reward Functions
F.2 The lowest income_bracket distribution tail is very important, focus heavily on the top 12.5%.
@@
F.3 Select solely those who speak Hindi to analyze how this would affect the calls made to our beneficiaries.
@@
F.4 Focus only on both the young and elderly.
@@
F.5 Weigth the lowest income_bracket groups, the lowest earners in the population.
@@
F.6 Combine the distributions of ’age’ and ’education’ to heavily give precedence to low income impoverished youth.
@@
F.7 Advantage those who prefer being called between 12:30pm-3:30pm and are registered at an NGO.
@@
F.8 Prefer the income_bracket bounds for the middle 40% of the population.
@@
F.9 Fixate on those Marathi-speakers with middle-aged folks.
@@
F.10 Give more attention for beneficiaries who likely work early in the morning and late at night.
@@
F.11 Infer technical challenges in reaching the phone that could indicate ’at-risk’ beneficiaries and give preference.

### 23 Feb 2024 v2

@@ 5.4 DLM Reflection Example
In Figure 3 we visualize state-feature distributions for two candidate rewards generated by the LLM. The LLM selected best index is highlighted with a green box. We show 4 (out of 14 total) relevant feature categories: age, income, education, and phone owners. We observe that the DLM selection is often able to reasonably interpret unambiguous commands, such as prioritizing older ages, and select the distribution that best achieves the given prompt. However, we note that for more ambiguous criteria, such as “technologically challenged”, the DLM selection choice is oftentimes more opaque; while the base reward for this prompt considers only income and phone ownership, additional features such as education level may also be considered for selection. In these cases, additional human feedback may be required to better disambiguate prompts for improved reflection. We provide full reward reflection logs for the below examples in Appendix F.
Figure 3: DLM reflection distribution and selection visualization. DLM uses outcome distributions from trained policy simulations to refine proposed reward functions. Above, we show the state-feature distributions for two rewards two prompts. DLM is able to reasonably interpret given prompts using distributions across feature groups; however, for more ambiguous prompts (“technologically challenged”), the selection criteria may become more opaque, potentially requiring additional human feedback.
@@ 5.5 Exceeding Base in Critical Low Income
We find that empirically for the “Critical Low Income” task the LLM is actually able to improve over the base reward. We find empirically that the DLM model often proposes a reward that is of much larger scaling than the original reward. For a representative example, we show a sample reward generated by the LLM for the Critical Low Income prompt:
@@
state + state * 10 * if_(agent_feats.no_or_lowest_income or agent_feats.second_lowest_income or agent_feats.third_lowest_income)
As our results in Figure 2 are evaluated in the low finite-sample regime, where there is not many gradient steps taken, sample efficiency becomes extremely important. In this case, because the reward function generated by the LLM gives an order of magnitude higher weighting to specific arms, it seems likely that the underlying policy gradient algorithm PPO gains sample efficiency in our setting. It is important to note that the improvement that the LLM yields over the base reward will shrink as the number of samples taken during training goes to infinity, as by definition the base reward defines the optimal policy. However, the highly shaped reward given by the LLM in this setting helps PPO converge faster for this task. It is also important to note that the base reward may not be the most sample efficient reward; other more complex rewards relying on correlated features or with different reward scaling will likely also gain efficiency. This also opens the possibility of achieving potential superhuman performance on our task in the future.
@@ 6 Conclusion
We propose a Decision-Language Model for resource allocation tasks in public health. The proposed model can take language prompts describing complex, potentially ambiguous policy outcome goals as input, and generate downstream control policies specialized for such goals. The proposed LLM-based approach can accurately capture the referenced prompt features and design effective reward functions in the form of code. Critically, our algorithms enable self-iteration on proposed reward functions in a fully automated fashion. By grounding LLM reward proposals with outcome distributions from RMAB simulations, we demonstrate the ability for the proposed DLM model to achieve near-human performance in public health.
@@
Future deployment of this technology could yield large societal benefit by improving public health outcomes and enabling more targeted policies. For our experiments, consent is obtained from every beneficiary enrolling in the ARMMAN’s mobile health program (see also Appendix A). However, constraining the deployed system to exhibiting only safe behavior is extremely important. One should take steps towards responsible deployment by performing human-in-the-loop verification before any learned policy allocation and implementing more traditional fallback systems not reliant on AI. We also address fairness in Appendix B.
APPENDIX:
A Decision-Language Model (DLM) for Dynamic Restless Multi-Armed Bandit Tasks in Public Health
@@ Appendix D Provided Tasks
We show provided tasks in Appendix E. We note that we evaluate on a wide variety of tasks within the real of resource allocation for maternal health. In particular, we test across feature characteristics such as age, income, spoken languages (an important characteristic indicating ethnicity in the Indian subcontienent), education, and other relevant features that may be considered in real-world allocation tasks. We provide these tasks directly as context to the LLM, in addition to relevant feature characteristics and their corresponding descriptions. Per seed, we fix the randomness in the sampled set of 48 arms from ARMMAN’s dataset, the deep RL neural network initialization, batch sampling from the dataset, and the sampling of states for each arm during evaluation. However please note that we cannot control the randomness of the Gemini LLM generation as we only have access through the Google API interface.
Appendix E Detailed Task Breakdown
Here we include the full prompt list for each task, base reward functions, as well as the numerical results given in Figure 2.
Label Task Command Base Function No Action Random Base Default DLM (No Reflection) DLM (Reflection)
@@ Appendix E Detailed Task Breakdown
Combine the distributions of ’age’ and ’education’ to heavily give precedence to low income impoverished youth.
@@
Infer technical challenges in reaching the phone that could indicate ’at-risk’ beneficiaries and give preference.
@@
Appendix F Sample Reward Reflection Outputs
@@
F.1 Task: Older Age Bias
@@ 35. Income bracket -1 (no income) - Binary
F.2 Task: Technologically Challenged
@@
Appendix G Generated Reward Functions
Here we give all generated reward functions seen in Appendix E from the LLM. We first give the task description, which is fed to the LLM, the corresponding ground truth base reward, and then the generated LLM rewards in order of seed. Note that we have processed the original raw indices into feature labels for interpretability.
G.1 Focus on the most elderly by age distribution.
@@ Appendix G Generated Reward Functions
G.2 The lowest income_bracket distribution tail is very important, focus heavily on the top 12.5%.
@@
G.3 Select solely those who speak Hindi to analyze how this would affect the calls made to our beneficiaries.
@@
G.4 Focus only on both the young and elderly.
@@
G.5 Weigth the lowest income_bracket groups, the lowest earners in the population.
@@
G.6 Combine the distributions of ’age’ and ’education’ to heavily give precedence to low income impoverished youth.
@@
G.7 Advantage those who prefer being called between 12:30pm-3:30pm and are registered at an NGO.
@@
G.8 Prefer the income_bracket bounds for the middle 40% of the population.
@@
G.9 Fixate on those Marathi-speakers with middle-aged folks.
@@
G.10 Give more attention for beneficiaries who likely work early in the morning and late at night.
@@
G.11 Infer technical challenges in reaching the phone that could indicate ’at-risk’ beneficiaries and give preference.